A copper cylinder is I initially at 20.0°C. At what temperature will it's volume be 0.150% larger than it is at 20.0°C


1
Expert's answer
2021-04-12T06:56:48-0400

The change of volume during the termal expansion is given as follows:


"\dfrac{\Delta V}{V} = \alpha (T_f -T_i)"

where "\Delta V/V = 0.15\% = 0.0015" is the relative volume change, "\alpha = 51\times 10^{-6}\space \degree C^{-1}" is the volumetric thermal expansion coefficient of copper, "T_i = 20.0\degree C" is the initial temperature, and "T_f" is the unknown final temperature. Expressing "T_f", find:


"T_f = T_i + \dfrac{1}{\alpha}\dfrac{\Delta V}{V}\\
T_f = 20\degree C + \dfrac{1}{51\times 10^{-6}\space \degree C^{-1}}\cdot 0.0015 \approx 49\degree C"

Answer. "49\degree C".